Jefferson Lab HALL A Logbook for February 02 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 634 02/28/02 21:40 khandaker FPP HV trip
  • 626 02/28/02 19:15 khandaker Q1 quenched
  • 622 02/28/02 18:25 atrig Trigger-Download
  • 607 02/28/02 16:17 R.A. Lindgre Shift summary
  • 4592 02/28/02 12:20 RAL shift half way summary
  • 4576 02/28/02 09:16 I.Rachek PID control is restored
  • 4573 02/28/02 08:41 RAL running status
  • 4568 02/28/02 08:06 Roedelbronn Owl shift summary
  • 4565 02/28/02 07:44 Roedelbronn Argon gas flow
  • 4555 02/28/02 05:41 Areg VDC efficiencies.
  • 4551 02/28/02 04:36 Roedelbronn beam is back, target control
  • 4549 02/28/02 03:33 Roedelbronn Target items
  • 4548 02/28/02 03:00 Roedelbronn problems
  • 4546 02/28/02 00:47 Roedelbronn ROC14 and shift turnover
  • 4544 02/28/02 00:01 Perdrisat 16 to 24 hour shift summary
  • Wednesday

  • 4541 02/27/02 23:46 Areg Run 3325-3329
  • 4515 02/27/02 17:19 R. Michaels spot++ explains itself now
  • 4509 02/27/02 16:35 perdrisa beam spot
  • 4505 02/27/02 15:57 Saha Shift Summary
  • 4493 02/27/02 13:49 R. Michaels synch check
  • 4453 02/27/02 08:09 Wendy Owl Shift Summary
  • 4419 02/27/02 00:45 Jack VDC high voltage
  • 4415 02/27/02 00:10 romanip HV frame(hv04) replaced
  • 4414 02/27/02 00:01 Solvignon Shift summary
  • Tuesday

  • 4408 02/26/02 22:46 Solvigno no beam
  • 4404 02/26/02 22:34 Solvigno Calorimeter HV
  • 4402 02/26/02 22:17 R. Michaels synch problems, decide to run unbuff
  • 4396 02/26/02 22:00 Jack VDC high voltage
  • 4388 02/26/02 20:52 R. Michaels prescale factor problem
  • 4380 02/26/02 20:07 Solvigno MCC
  • 4379 02/26/02 20:05 Solvigno CODA problem.
  • 4377 02/26/02 19:25 Solvigno VDC HV
  • 4375 02/26/02 19:21 Jack VDC power supplies
  • 4360 02/26/02 17:20 x. jiang record of FPP chamber.
  • 4354 02/26/02 16:17 x. jiang no FPP trips during the day.
  • 4352 02/26/02 13:21 R. Michaels R-arm DAQ
  • 4348 02/26/02 13:14 romanip Veto HVframes are removed
  • 4346 02/26/02 11:22 Saha Shift Summary
  • 4344 02/26/02 11:07 Saha Fasrt Feedback running off
  • 4321 02/26/02 07:58 Kumbartzki shift summary
  • 4263 02/26/02 00:00 Punjabi Shift Summary
  • Monday

  • 4259 02/25/02 23:45 tchang coincidence window and threshold change
  • 4195 02/25/02 16:03 LeRose Shift Summary
  • 4189 02/25/02 15:34 Kumbartzki Cryo target
  • 4176 02/25/02 14:37 Higinbotham Machine Set Energy
  • 4171 02/25/02 14:25 R. Michaels DAQ problems, fixed, we think
  • 4167 02/25/02 13:23 LeRose Hall Access
  • 4139 02/25/02 10:11 tchang beam current vs T1 counting rate
  • 4137 02/25/02 09:31 LeRose More Magnet Trips
  • 4136 02/25/02 09:07 LeRose Left Dipole Trip
  • 4127 02/25/02 08:00 C. Hyde-Wrig Shift Summary
  • 4116 02/25/02 07:09 Bosted TS2 died
  • 4113 02/25/02 06:45 kramerk FPP HV tripped again
  • 4090 02/25/02 04:32 Hyde-Wright VDC
  • 4088 02/25/02 04:19 kramerk Calorimeter HV tripped
  • 4077 02/25/02 03:01 kramerk TS2 rebooted
  • 4072 02/25/02 02:26 C.Hyde- Wrig VDC
  • 4065 02/25/02 00:58 C. Hyde-Wrig argon VDC flow
  • 4062 02/25/02 00:44 kramerk DAQ problem
  • 4055 02/25/02 00:09 Kumbartzki shift summary
  • Sunday

  • 4045 02/24/02 23:07 Jack Flammable gas alarms disabled
  • 3983 02/24/02 16:03 bertin run sumary
  • 3953 02/24/02 13:03 geraud Oscillations reduce
  • 3950 02/24/02 12:56 geraud target HPH oscillations
  • 3947 02/24/02 12:44 geraud Target High Power Heater
  • 3945 02/24/02 12:32 geraud scaler fixed
  • 3932 02/24/02 11:21 geraud scalers crash
  • 3900 02/24/02 08:00 L.Pentchev owl shift summary
  • 3883 02/24/02 06:02 Wendy target ioc reboot
  • 3831 02/24/02 00:11 Z.-E Meziani shift summary
  • Saturday

  • 3815 02/23/02 22:35 Z.-E. Mezian target, heat load
  • 3811 02/23/02 22:17 Areg problematic run
  • 3758 02/23/02 16:05 C. C. Chang Shift Summary
  • 3737 02/23/02 14:03 C. C. Chang Daq
  • 3722 02/23/02 12:42 C. C. Chang FPP Gas Flow
  • 3720 02/23/02 12:36 C. C. Chang VDC Low Voltage Control
  • 3704 02/23/02 10:56 tchang roc4 swapped
  • 3701 02/23/02 09:26 X. Zheng target ioc reboot
  • 3699 02/23/02 09:19 x. jiang the right arm door won't open.
  • 3698 02/23/02 09:17 x.jiang fpp gas flow.
  • 3689 02/23/02 07:56 L.Pentchev owl shift summary
  • 3686 02/23/02 07:48 L.Pentchev Ar gas flow problem
  • 3683 02/23/02 07:42 Jack Gas mixing system
  • 3674 02/23/02 06:48 David & Areg gas problems
  • 3616 02/23/02 00:43 perdrisat shift summary
  • Friday

  • 3587 02/22/02 18:59 tchang deadtime vs (un)buffered mode
  • 3557 02/22/02 16:03 LeRose shift summary
  • 3548 02/22/02 14:52 LeRose Magnet trips
  • 3547 02/22/02 14:14 LeRose Magnet Trip (Q1 too!)
  • 3546 02/22/02 14:04 LeRose Left Dipole Trip
  • 3545 02/22/02 14:03 LeRose Left Dipole Trip
  • 3544 02/22/02 13:49 Reitz EP measurement
  • 3543 02/22/02 13:12 atrig Trigger-Download
  • 3542 02/22/02 13:10 R. Michaels DAQ problems
  • 3530 02/22/02 11:44 LeRose Back to 3Ee Kinematics
  • 3515 02/22/02 08:45 LeRose Magnet Screen
  • 3513 02/22/02 08:03 Roedelbronn Owl shift summary
  • 3502 02/22/02 04:22 R.Pomatsalyu Moller measurement
  • 3501 02/22/02 02:41 tchang, Ed ROC4 problem persists
  • 3491 02/22/02 00:18 hyde Swing shift summary
  • Thursday

  • 3417 02/21/02 16:01 Saha Shift Summary
  • 3416 02/21/02 16:00 tchang coincidence window changed to 48 ns
  • 0 02/21/02 16:00 no_name no_key
  • 83400 02/21/02 14:44 X. Jiang to reset FPP HV channel 8.1.22
  • 83389 02/21/02 14:00 Saha calorimeter HV tripped
  • 83387 02/21/02 13:51 Saha FPP HV trip
  • 83357 02/21/02 08:10 I.Rachek ROC14 problems
  • 83355 02/21/02 07:58 X. Zheng shift summary
  • 83329 02/21/02 03:52 I.Rachek TS1 rebooted
  • 83322 02/21/02 02:36 tchang ROC4 in pedestal surpression mode
  • 83313 02/21/02 01:08 I.Rachek TS2 rebooted
  • 83310 02/21/02 00:33 X. Jiang Mott measurements at the injector
  • 83307 02/21/02 00:03 V. Sulkosky Swing shift summary
  • Wednesday

  • 83275 02/20/02 16:54 Jack Flammable gas sensors
  • 83274 02/20/02 16:53 Jack Gas flow
  • 83272 02/20/02 16:20 atrig Trigger-Download
  • 83270 02/20/02 12:52 tchang coincidence window adjusted
  • 83268 02/20/02 11:06 x. jiang the gas flow on the left arm
  • 83265 02/20/02 10:57 X. Jiang left arm gas flow adjusted to the detectors.
  • 83261 02/20/02 07:16 Wendy Owl Shift Summary
  • 83234 02/20/02 03:50 I.Rachek HV trip in calorimeter
  • Tuesday

  • 83217 02/19/02 23:59 V. Sulkosky Swing Shift summary
  • 83206 02/19/02 16:55 atrig Trigger-Download
  • 83201 02/19/02 15:50 Seonho Choi Shift Summary (Feb. 19 day)
  • 83191 02/19/02 11:17 Seonho Choi Incorrect angle for runs 2680, 2681, 2682
  • 83176 02/19/02 08:07 Wendy Owl Shift Summary
  • 83172 02/19/02 07:37 atrig Trigger-Download
  • 83171 02/19/02 06:55 Wendy Shift Update
  • 83153 02/19/02 02:52 Wendy Current
  • 83148 02/19/02 01:38 Wendy Shift updates
  • 83140 02/19/02 00:06 kramerk Optics Run done with raster on
  • 83139 02/19/02 00:04 atrig Trigger-Download
  • Monday

  • 83137 02/18/02 23:56 kramerk Moved Calorimeter to 3C config.
  • 83136 02/18/02 23:56 kramerk Shift Summary
  • 83128 02/18/02 22:51 kramerk Moved to Empty target and turned magnet off
  • 83127 02/18/02 22:40 kramerk Paged Rusty for Calorimeter Move
  • 83120 02/18/02 22:06 kramerk Moved to 3Df*
  • 83107 02/18/02 20:12 Areg target move
  • 83100 02/18/02 19:29 kramerk Another DAQ crash
  • 83096 02/18/02 19:16 Eugene More on the previous Moller measurement
  • 83092 02/18/02 18:19 kramerk Switched to 3De* kinematic
  • 83091 02/18/02 18:17 romanip HV channel trip hv01:L15.0
  • 83089 02/18/02 18:06 kramerk Requested 2uA
  • 83084 02/18/02 17:47 kramerk More DAQ problems
  • 83083 02/18/02 17:29 kramerk Calorimeter problems
  • 83077 02/18/02 16:40 kramerk FPP HV tripped again
  • 83076 02/18/02 16:34 kramerk HV trip in FPP
  • 83073 02/18/02 16:19 kramerk ROC14 crashed - reset and reboot
  • 83070 02/18/02 16:05 Rikki day shift summary
  • 83025 02/18/02 08:01 Wendy Owl Shift Summary
  • Sunday

  • 82989 02/17/02 23:55 Lingyan Swing shift summary
  • 82986 02/17/02 22:44 Eugene Moller results
  • 82983 02/17/02 22:24 Yuri Calorimeter HVs were changed
  • 82978 02/17/02 21:22 Lingyan calo HV
  • 82976 02/17/02 21:18 X. Jiang To reset FPP high voltage channel through LeCroy HV control GUI
  • 82972 02/17/02 20:47 Lingyan Status
  • 82970 02/17/02 20:38 atrig Trigger-Download
  • 82958 02/17/02 17:56 X. Zheng missing plot
  • 82957 02/17/02 17:53 X. Zheng left arm helicity check
  • 82954 02/17/02 17:35 tchang wein angle and 1/2 wave plate info
  • 82933 02/17/02 08:12 Roedelbronn Owl shift summary
  • 82908 02/17/02 02:43 atrig Trigger-Download
  • 82907 02/17/02 02:42 atrig Trigger-Download
  • 82891 02/17/02 00:04 C. C. Chang Run Summary
  • Saturday

  • 82879 02/16/02 22:36 Lingyan Target ready
  • 82874 02/16/02 19:03 Lingyan Target update
  • 82872 02/16/02 02:15 Run coordina Status and upcoming shifts
  • 82871 02/16/02 02:11 I.Rachek New pulser introduced to measure trigger deadtime
  • 82870 02/16/02 02:03 I.Rachek Oscillation in SUM32 #54 removed.
  • 82869 02/16/02 01:18 ddutta Shift Summary
  • 82868 02/16/02 00:05 Chang Run Summary
  • Friday

  • 82860 02/15/02 23:28 Saha Harp Scans
  • 82857 02/15/02 23:14 C. C. Chang Harp scan
  • 82846 02/15/02 21:06 Punjabi magnet settings
  • 82834 02/15/02 19:10 jpchen cryotarget condition
  • 82819 02/15/02 17:40 atrig Trigger-Download
  • 82818 02/15/02 17:39 atrig Trigger-Download
  • 82817 02/15/02 17:38 atrig Trigger-Download
  • 82816 02/15/02 17:37 atrig Trigger-Download
  • 82804 02/15/02 16:32 Adaq Harp Scan
  • 82802 02/15/02 13:57 Alan Nathan calorimeter absorber
  • 82801 02/15/02 13:44 atrig Trigger-Download
  • 82799 02/15/02 13:10 LeRose Spectrometer set
  • 82798 02/15/02 12:44 gilman cryotarget
  • 82797 02/15/02 12:35 M Jones revised estimate of calibration time for Ac measurement
  • 82792 02/15/02 08:24 gilman cryotarget work
  • 82790 02/15/02 06:45 dutta shift summary
  • 82787 02/15/02 03:24 Meekins motion of the cryotarget during cooldown/vacuum
  • 82785 02/15/02 00:00 E.Chudakov Hall status and the coming program
  • Thursday

  • 82781 02/14/02 17:55 M Jones Time estimate for Ac calibration from ep elastics
  • 82771 02/14/02 09:16 X. Jiang FPP threshold and low voltage screen
  • 82770 02/14/02 09:14 X. Jiang FPP high voltage channels
  • 82769 02/14/02 00:04 Jack VDC gas tubing
  • Wednesday

  • 82757 02/13/02 13:56 Reitz T4 on Scaler
  • 82752 02/13/02 09:16 Higinbotham Web Server Updated
  • Tuesday

  • 82746 02/12/02 20:32 atrig Trigger-Download
  • 82743 02/12/02 20:26 atrig Trigger-Download
  • 82740 02/12/02 20:19 atrig Trigger-Download
  • 82737 02/12/02 20:14 atrig Trigger-Download
  • 82734 02/12/02 20:07 atrig Trigger-Download
  • 82731 02/12/02 19:58 atrig Trigger-Download
  • 82728 02/12/02 19:49 atrig Trigger-Download
  • 82725 02/12/02 19:43 atrig Trigger-Download
  • 82722 02/12/02 19:36 atrig Trigger-Download
  • 82719 02/12/02 19:29 atrig Trigger-Download
  • 82716 02/12/02 19:24 atrig Trigger-Download
  • 82713 02/12/02 19:18 atrig Trigger-Download
  • 82710 02/12/02 19:04 atrig Trigger-Download
  • 82707 02/12/02 18:58 atrig Trigger-Download
  • 82706 02/12/02 18:47 atrig Trigger-Download
  • 82703 02/12/02 18:41 atrig Trigger-Download
  • 82700 02/12/02 18:35 atrig Trigger-Download
  • 82697 02/12/02 18:28 atrig Trigger-Download
  • 82696 02/12/02 18:26 atrig Trigger-Download
  • 82692 02/12/02 18:15 atrig Trigger-Download
  • 82685 02/12/02 16:51 Igor&Yuri Fast cable for T4 installed
  • 82679 02/12/02 14:29 Yuri cosmic trigger paddles are on.
  • 82676 02/12/02 08:42 Mike & David LH2 Target
  • 82671 02/12/02 06:03 dutta shift summary
  • 82665 02/12/02 04:02 Roedelbronn veto ON
  • 82662 02/12/02 03:22 atrig Trigger-Download
  • 82661 02/12/02 03:22 Roedelbronn Problems with HRS-L diplole
  • 82659 02/12/02 00:43 Roedelbronn RCS HV trip
  • 82655 02/12/02 00:03 Guy Ron Shift Summary
  • Monday

  • 82643 02/11/02 20:27 Guy Ron Spectrometer set-up
  • 82638 02/11/02 19:00 Reitz EP Energy
  • 82635 02/11/02 18:26 tchang FPP CH2 position moved
  • 82633 02/11/02 18:21 Guy Ron HV
  • 82626 02/11/02 15:55 dhamilton Shift Summary
  • 82625 02/11/02 15:30 jpchen scattering chamber vacuum readback
  • 82624 02/11/02 15:26 dhamilton RCS HV Trip
  • 82621 02/11/02 14:53 R. Michaels T4 in scalers
  • 82620 02/11/02 14:51 dhamilton RCS Magnet On
  • 82618 02/11/02 14:43 Saha Hall A Arc Energy Measurement
  • 82617 02/11/02 11:05 jpchen target moved to optics target
  • 82611 02/11/02 09:56 dhamilton ROC4 & 14 Problems
  • 82609 02/11/02 09:26 dhamilton RCS Magnet On
  • 82603 02/11/02 08:00 X. Zheng shift summary
  • 82595 02/11/02 05:31 X. Zheng beam current 20uA
  • 82575 02/11/02 02:05 X. Zheng Beam current is 30 uA
  • 82573 02/11/02 01:42 tchang TS2 and ROC4 hub replaced
  • 82572 02/11/02 01:24 adaq correction to run numbers
  • 82571 02/11/02 01:17 X.Zheng controlled access
  • 82567 02/11/02 00:04 E.Chudakov Shift summary
  • Sunday

  • 82558 02/10/02 22:56 Eugene Beam current is 40uA
  • 82554 02/10/02 22:11 Eugene Beam current reduced to 35uA
  • 82545 02/10/02 20:17 Eugene RCS magnet off
  • 82544 02/10/02 20:16 Eugene T4 delay
  • 82539 02/10/02 19:27 Eugene CODA pauses
  • 82535 02/10/02 19:15 Eugene Data decoding problem
  • 82513 02/10/02 16:04 B. Vlahovic Shift summary
  • 82505 02/10/02 15:09 atrig Trigger-Download
  • 82493 02/10/02 11:50 Eugene RCS HV wrong setting for Jan,9 - Jan,10
  • 82491 02/10/02 11:31 Yuri target status
  • 82483 02/10/02 08:03 X. Zheng shift summary
  • 82471 02/10/02 05:20 X. Zheng Reduce beam current from 25uA to 20uA
  • 82461 02/10/02 03:31 I.Rachek Coolant flow reduced
  • 82446 02/10/02 00:53 R. Michaels T4 time offset ok (but now kine dependent)
  • 82445 02/10/02 00:48 atrig Trigger-Download
  • 82442 02/10/02 00:06 Punjabi Shift Summary
  • Saturday

  • 82435 02/09/02 23:40 atrig Trigger-Download
  • 82428 02/09/02 22:32 R. Michaels T4 time offset
  • 82420 02/09/02 21:14 bogdan For attention of Bob Michaels
  • 82415 02/09/02 20:00 romanip HV channel trips and HV frame problems
  • 82413 02/09/02 17:46 atrig Trigger-Download
  • 82412 02/09/02 17:44 atrig Trigger-Download
  • 82400 02/09/02 16:00 B. Vlahovic The shift summary
  • 82386 02/09/02 14:41 vladimir nel HV status of calorimeter
  • 82380 02/09/02 14:22 vladimir nel target status
  • 82361 02/09/02 11:34 vladimir nel target status
  • 82360 02/09/02 11:27 Vladimir Nel Target status
  • 82353 02/09/02 08:19 dhamilton FPP back on
  • 82350 02/09/02 08:07 V. Sulkosky Owl Shift summary
  • 82345 02/09/02 07:26 V. Sulkosky Magnet off
  • 82343 02/09/02 07:25 I.Rachek veto VH16 high voltage problem
  • 82322 02/09/02 03:22 V. Sulkosky Argon gas pressure low
  • 82321 02/09/02 03:13 dhamilton FPP off
  • 82318 02/09/02 02:11 V. Sulkosky Move to offendpt Kin 5D
  • 82305 02/09/02 00:09 Vladimir Nel Swing shift summary
  • Friday

  • 82294 02/08/02 22:19 atrig Trigger-Download
  • 82293 02/08/02 22:18 atrig Trigger-Download
  • 82292 02/08/02 22:15 x. Jiang reset iocha48, big box control ioc.
  • 82254 02/08/02 16:01 gilman shift summary
  • 82214 02/08/02 11:14 Saha Arc Integral Measurement
  • 82192 02/08/02 08:02 E.Chudakov Shift summary
  • 82186 02/08/02 07:19 Eugene RCS hub problem
  • 82185 02/08/02 07:18 Eugene Access
  • 82183 02/08/02 06:25 Eugene TS2, ROC4 got stuck
  • 82159 02/08/02 03:25 E.Chudakov TS2 crash
  • 82144 02/08/02 00:58 x. jiang finished kin-5G
  • 82143 02/08/02 00:56 atrig Trigger-Download
  • 82141 02/08/02 00:03 LIndgren Shift Summary
  • Thursday

  • 82112 02/07/02 16:02 gilman shift summary II
  • 82111 02/07/02 15:59 gilman shift summary
  • 82106 02/07/02 15:20 gilman calo threshold
  • 82105 02/07/02 15:20 Eugene RCS HV spare crate
  • 82089 02/07/02 13:45 gilman for w calo peds
  • 82080 02/07/02 13:02 gilman calo threshold reset
  • 82074 02/07/02 12:15 gilman reset calorimeter threshold
  • 82064 02/07/02 08:23 ddutta shift summary
  • 82059 02/07/02 06:34 Rachek Calorimeter HV tripped
  • 82051 02/07/02 04:16 ddutta DAQ crashes
  • 82039 02/07/02 00:30 C. Hyde-Wrig Beam Energy
  • 82034 02/07/02 00:01 Punjabi Shift summary
  • Wednesday

  • 82030 02/06/02 22:32 atrig Trigger-Download
  • 82029 02/06/02 22:25 atrig Trigger-Download
  • 82009 02/06/02 18:55 aregjan vdc multiplicities
  • 81997 02/06/02 16:19 C.Hyde-Wrigh Beam Time Projections
  • 81992 02/06/02 16:03 jpchen day shift summary
  • 81991 02/06/02 15:59 jpchen LH2 pressure readings increasing
  • 81970 02/06/02 12:13 Saha BCM Calibration - 0L02/FC2
  • 81946 02/06/02 07:59 Annand Owl Shift Summary
  • 81931 02/06/02 06:17 Annand LH2 + Radiator Target In
  • 81929 02/06/02 05:41 Annand TS1 crash
  • 81923 02/06/02 04:13 Annand HRSL setting
  • 81919 02/06/02 02:55 Annand RCS Magnet ON
  • 81900 02/06/02 00:05 markowitz shift summary
  • Tuesday

  • 81898 02/05/02 21:13 Chudakov HV change for RCS
  • 81895 02/05/02 20:41 markowit trigsetup again
  • 81892 02/05/02 20:34 atrig Trigger-Download
  • 81891 02/05/02 20:33 atrig Trigger-Download
  • 81886 02/05/02 20:18 Yuri replacement of LG phototube (channel #263)
  • 81887 02/05/02 20:15 markowitz trigsetup
  • 81885 02/05/02 20:09 atrig Trigger-Download
  • 81884 02/05/02 20:03 *C.*H.-*W. 0L02
  • 81883 02/05/02 19:51 C. Hyde-Wrig BCM
  • 81881 02/05/02 18:27 Jack VDC gas flow
  • 81876 02/05/02 17:33 Hyde-Wright OL02 vs FC2 Calibration
  • 81866 02/05/02 16:48 E.Chudakov Shift summary
  • 81864 02/05/02 15:19 L.Pentchev IOC reboot
  • 81863 02/05/02 14:49 Eugene LH target: IOC reboot
  • 81860 02/05/02 14:24 R. Michaels how to measure charge asy
  • 81856 02/05/02 12:44 Eugene Moller results
  • 81851 02/05/02 07:59 Annand Owl Shift Summary
  • 81847 02/05/02 07:07 Annand Beam OFF
  • 81844 02/05/02 06:32 Annand Optics Target
  • 81837 02/05/02 05:28 Annand RCS Magnet OFF
  • 81832 02/05/02 04:29 Annand QE measurement
  • 81814 02/05/02 01:10 Annand Calorimeter TS2 Reset
  • 81813 02/05/02 00:49 Annand HV, CODA crash
  • 81812 02/05/02 00:38 Annand Calorimeter High Voltage
  • Monday

  • 81805 02/04/02 23:57 markowitz shift summary
  • 81779 02/04/02 19:29 atrig Trigger-Download
  • 81775 02/04/02 19:08 areg magnet ON/OFF
  • 81753 02/04/02 16:00 LeRose Shift Summary
  • 81752 02/04/02 15:55 atrig Trigger-Download
  • 81751 02/04/02 14:31 LeRose Kinematics change (to 5c)
  • 81748 02/04/02 14:02 LeRose Optics check
  • 81743 02/04/02 13:19 LeRose Back to 10 microamps
  • 81739 02/04/02 12:19 LeRose Beam I @ 2.5
  • 81737 02/04/02 11:56 LeRose Beam Current to 3
  • 81733 02/04/02 11:09 LeRose correction
  • 81730 02/04/02 10:50 LeRose Beam position Adjustment
  • 81727 02/04/02 10:22 LeRose Reduced Thresholds
  • 81721 02/04/02 08:03 R. Michaels owl shift summary
  • 81707 02/04/02 04:57 R. Michaels CODA problems (fixed)
  • 81698 02/04/02 04:02 R. Michaels trigger latch pattern
  • 81691 02/04/02 03:06 R. Michaels start kinematics 5A
  • 81685 02/04/02 02:14 atrig Trigger-Download
  • 81684 02/04/02 02:03 R. Michaels VDC efficency recovery timescale
  • 81683 02/04/02 01:56 Rachek Set Optics Target
  • 81677 02/04/02 00:29 R. Michaels VDC efficiency
  • Sunday

  • 81676 02/03/02 23:53 Markowitz shift summary
  • 81669 02/03/02 20:56 Jack Gas system status
  • 81668 02/03/02 19:37 Jack Argon contamination time
  • 81667 02/03/02 19:31 Jack Gas mixing system gas
  • 81662 02/03/02 18:29 Jack Gas mixing system woes
  • 81660 02/03/02 17:51 adaq right arm HVsetting
  • 81659 02/03/02 17:25 tchang RCS trigger pattern latch connected
  • 81656 02/03/02 17:01 Eugene DVCS tests with RCS calorimeter
  • 81655 02/03/02 16:42 Adaq angle to 38.47
  • 81651 02/03/02 16:15 kramerk Shift Summary
  • 81650 02/03/02 15:51 gunning raster
  • 81647 02/03/02 14:32 C. Hyde-Wrig DVCS test
  • 81646 02/03/02 14:10 C.Hyde-Wrigh VDC
  • 81643 02/03/02 13:57 kramerk Hall A in Controlled Access
  • 81639 02/03/02 13:41 C.Hyde-Wrigh Status- 24 hours
  • 81638 02/03/02 13:36 kramerk Gas Shed Problem - Pressed Low Pressure Override
  • 81636 02/03/02 13:27 dhamilton FPP off
  • 81634 02/03/02 13:22 kramerk Gas Alarm went off
  • 81631 02/03/02 13:10 Hyde-Wright raster spot++
  • 81630 02/03/02 13:07 dhamilton espace problems
  • 81621 02/03/02 12:07 kramerk MCC sending production beam
  • 81620 02/03/02 11:45 kramerk MCC sending pulse beam
  • 81619 02/03/02 11:27 kramerk MCC delivering 1 uA
  • 81616 02/03/02 08:33 kramerk MCC sending some beam
  • 81614 02/03/02 08:02 R. Michaels owl summary
  • 81607 02/03/02 03:45 atrig Trigger-Download
  • 81606 02/03/02 03:38 Eugene Access to reset HV06
  • Saturday

  • 81593 02/02/02 23:56 Markowitz shift summary
  • 81591 02/02/02 21:46 atrig Trigger-Download
  • 81590 02/02/02 21:42 atrig Trigger-Download
  • 81579 02/02/02 17:58 Markowitz LH2 sensor
  • 81578 02/02/02 16:48 Adaq runs 1991 and 1992 incorrect angle
  • 81573 02/02/02 16:05 kramerk Shift Summary
  • 81570 02/02/02 15:59 kramerk Moved spectrometer 21.829
  • 81568 02/02/02 15:40 kramerk Moved spectrometer to 21.297
  • 81564 02/02/02 15:26 kramerk Moved spectrometer to 20.764
  • 81559 02/02/02 14:47 kramerk Spectrometer moved to 20.232
  • 81552 02/02/02 14:00 kramerk Target moved to LH2 no rad.
  • 81551 02/02/02 13:59 kramerk Harp Scan Raster On
  • 81550 02/02/02 13:58 dhamilton target move 15rad
  • 81547 02/02/02 13:38 karmerk MCC investigating some problems
  • 81546 02/02/02 13:26 kramerk OTR info - Raster On
  • 81544 02/02/02 13:12 Hyde-Wright Raster at 5 pass
  • 81542 02/02/02 13:00 C. Hyde-Wrig Raster
  • 81536 02/02/02 12:43 kramerk Harp Scan at 5.7 GeV #1
  • 81532 02/02/02 12:28 C. Hyde-Wrig raster
  • 81527 02/02/02 11:31 Albert Post-Radiator geometry
  • 81522 02/02/02 08:21 atrig Trigger-Download
  • 81520 02/02/02 08:04 Roedelbronn Shift summary
  • 81516 02/02/02 07:04 atrig Trigger-Download
  • Friday

  • 81502 02/01/02 17:52 markowitz mcc tuning pulse length
  • 81495 02/01/02 17:10 markowitz spectrometer setting change
  • 81485 02/01/02 15:56 Kumbartzki day shift summary
  • 81482 02/01/02 15:40 Kumbartzki FPP
  • 81481 02/01/02 15:26 Rachek Cables Damage Report
  • 81479 02/01/02 14:54 C. Hyde-Wrig Configuration Change
  • 81478 02/01/02 13:01 C. Hyde-Wrig HRS-L
  • 81477 02/01/02 12:56 C. Hyde-Wrig Access
  • 81473 02/01/02 12:05 Kumbartzki general
  • 81470 02/01/02 11:40 Jack Gas system trip
  • 81469 02/01/02 11:37 Charles Hyde HRS Motion
  • 81464 02/01/02 09:45 Kumbartzki FPP
  • 81463 02/01/02 09:28 todor Gas Alarm
  • 81454 02/01/02 08:00 Roedelbronn Shift summary
  • 81437 02/01/02 03:44 atrig Trigger-Download
  • 81416 02/01/02 00:03 nsavvinv swing shift summary